OpenStudy (anonymous):

What is the length of a rectangle that has a width of 9.2 centimeters and an area of 115.92 square centimeters?

OpenStudy (anonymous):

1066.46 cm 12.6 cm 102.6 cm 15 cm

OpenStudy (anonymous):

1.10*X=40 2. Divide each side by 10: X=4 centimeters 3. Check your answer 10*4=40

OpenStudy (lncognlto):

A = l*w 115.92 = l*9.2 therefore l = 115.92/9.2

OpenStudy (petiteme):

Area = length * width so if length is unknown; area/width = length length=115.92 divided by 9.2 :)